A computer prints out three digits chosen at random from 0 to 9, inclusive. Find the probability that:
a) All three digits are different.
b) The third digit differs from the first two digits.
Thank yoy very much for your answer!

Answer by edjones(7569)

a) 10/10 * 9/10 * 8/10 = 720/1000 = 18/25
.
b)
((8/10 * 9) + 9/10)/10 Because of a 1/10 chance both of the first 2 digits are the same.
= (72/10 + 9/10)/10
= 81/100
.
Ed